OpenStudy (anonymous):

Multiply. (7b+7)^2 I get confused,do I use FOIL in this?

OpenStudy (kuoministers):

you dont have too. use a^2 + 2ab + b^2

hero (hero):

I prefer distributive rule: (7b + 7)(7b + 7) = 7b(7b + 7) + 7(7b + 7) = 49b^2 + 49b + 49b + 49 = 49b^2 + 98b + 49

hero (hero):

You know, distributive rule is an actual mathematical property. FOIL is just an acroymn someone made up.

OpenStudy (anonymous):

Ah, ok I see now. I didn't know that,but it helps me sadly to think of First, Outer, Inner and Last

hero (hero):

FOIL is something kids in high school learn.

OpenStudy (anonymous):

distibutive is the easiest way of doing this. I learned distributive last year. helped alot.

OpenStudy (anonymous):

Lol well they're teaching it in our college. Either way thank you so much for your help